Leggett and Platt is a world leader in the manufacturing of a broad line of components for the home, office and institutional furnishings and has a plant located in Simpsonville, KY. The following financial information of this plant is modified (for obvious reasons) in order to illustrate some of TOC concepts related to financial and operational measures.

Partial information from Income Statement:

Sales Revenue                                                                                                                     $50,000,000

Cost of Goods Sold (COGS)                                                                                              $35,000,000

(Direct Material = 40%, Labor and overheads = 60%)

Selling & Marketing Expenses                                                                                         $    500,000

General Administrative Expenses                                                                                    $ 2,000,000

Partial information from Balance Sheet:

Inventory (Direct Material, Work-in-progress and Finished Goods)                         $10,000,000

Other Assets (Machinery, Computers, Buildings etc.)                                                 $25,000,000

Notes: Assume that (i) Direct Material portion of Cost of Goods Sold is about 40%, and (ii) Direct material portion of Inventory as appeared in Balance Sheet is about 50%.

Based on the above limited financial information gathered from Balance Sheet and Income Statement:

1. Calculate following financial and operational measures as we have discussed in the class:

(i) Throughout,

(ii) Inventory,

(iii) Operating Expenses,

(iv) Net Profit, and

(v) Return on Inventory

1.For TOC Mindset and Measures​ homework, what is the throughput for Leggett and​ Platt?

A.$​40,000,000

B.$​15,000,000

C.$36,000,000

D.​$29,000,000

For TOC Mindset and Measurement​ homework, what is the TOC Inventory for Leggett and​ Platt?

A.$39,000,000

B.​$30,000,000

C.​$24,000,000

D.​$38,000,000

For TOC Mindset and Measures homework​ problem, what are the Oerating Epenses for Leggett and​ Platt?

A.​$23,500,000

B.$21,000,000

C.$15,000,000

D.​$21,500,000

For TOC Mindset and Measures homework​ problem, what is the​ Return-On-Investment (ROI) for Leggett and​ Platt?

A.50%

B.55%

C.41.667%

D.​32.895%
Answer #1

(i)

Given that,

Sales Revenue = 50000000

Cost of good Sold = 35000000

Direct Material = 40% of Cost of Goods Sold = 35000000 * 40% = 14000000

Throughput = Sales Revenue - Direct Material = 50000000 - 14000000 = 36000000

(ii)

Total Inventory = 10000000

Direct Material Inventory = 50% * 10000000 = 5000000

iii)

Operating Expense = All expenses excluding Direct Material

Operating Expense = 21000000 + 500000 + 2000000 = 23500000

iv)

Net Profit = throughput - operating expense = 36000000 - 23500000 = 12500000

v) Return of Investment (Inventory) = Net Profit/ Direct Material Cost = 12500000/14000000 = 89.29%